Le Terroir Winery is considering an expansion project to produce fine wines. The trial expansion will last for 4 years. It will require a capital investment of $1,000,000, and additional net operating working capital of $150,000. The winery expects to sell 10,000 bottles of wine each year at $98/bottle. The variable cost is $60/bottle, and total fixed costs are $200,000. Each of these are expected to be constant over the 4 year life of the project.

a. Assuming a tax rate of 40%, WACC of 10%, straight-line depreciation, and a salvage value of $50,000, evaluate the NPV of the project.

b. Perform sensitivity analysis to unit price—let there be a 40% chance price is $20, and a 60% that price is $150. Determine the new NPV based on these possible outcomes. c. If the wine is successful and price is $150, there will be spillover effects to the company’s other wines. This will lead to company growth of approximately $1,000,000 in value at the end of the project. Estimate the new NPV of the project.
Answer #1

a. NPV of the project:

Computation of annual cash inflow:

Particulars Computation Amount ($)
Sales 10,000 bottles x $98 980,000
Less: Variable Costs 10,000 bottles x $60 600,000
Contribution Margin 380,000
Less: Fixed Costs 200,000
Less: Depreciation (1,000,000 - 50,000) / 4 237,500
Profit / (loss) before tax (57,500)
Less: Tax at 40% (Savings) (57,500) x 40% (23,000)
Profit after Tax -57,500 - (-23,000) (34,500)
Add: Depreciation 237,500
Annual Cash inflow 203,000

* To arrive at annual cash flows, Tax effect needs to be considered, for which first depreciation needs to be reduced to arrive at accounting profit / (loss). When there is a loss, it means that there is a tax benefit. Since Depreciation is a non-cash item, the same is added back to Profit after Tax in order to arrive at Annual Cash flow.

In the 4th year, there is additional cash inflow in the form of salvage value amounting to $50,000.

Net present value (NPV) is the present value of all cash flows. In other words, it is the difference between the present value of cash inflows and present value of cash outflows. It takes into account the time value of money.

Weighted average cost of capital (WACC) is used to arrive at present value of cash flows.

Present Value of cash flow = Cash flow / (1 + rate)year

However, when the cash flows are uniform, then Present value of cash flows can be arrived by

Present Value of cash flows = Annual Cash flow   x    Annual annuity factor

Annuity factor of WACC 10% for 5 years = [1 - 1 / (1+0.10)5] / 0.10 = 3.7908

Present Value of Annual Cash flows = $203,000 x 3.7908 = $769,532.40

Present Value of Salvage Value = $50,000 / (1+0.10)5 = $31,046

Total Present value of cash inflows = $800,578.40

Present value of cash outflows = initial investment = $1,000,000

Net Present Value = 800,578.40 - 1,000,000 = $ (199,421.60)

Since the NPV is negative, the project is not feasible at price of $98 per bottle.

b. Sensitivity analysis:

The Pessimistic sale price is$20 at 40% chance and optimistic sale price is $150 at 60% chance.

The expected sale price is (20 x 40%) + (150 x 60%) = $98 per bottle.

NPV when expected sale price is $98/bottle is $ (199,421.60) as arrived in (a) above.

c. NPV when price is $150:

Computation of annual cash inflow:

Particulars Computation Amount ($)
Sales 10,000 bottles x $150 1,500,000
Less: Variable Costs 10,000 bottles x $60 600,000
Contribution Margin 900,000
Less: Fixed Costs 200,000
Less: Depreciation (1,000,000 - 50,000) / 4 237,500
Profit / (loss) before tax 462,500
Less: Tax at 40% 462,500 x 40% 185,000
Profit after Tax 462,500 - 185,000 277,500
Add: Depreciation 237,500
Annual Cash inflow 515,000

In the 4th year, there is additional cash inflow in the form of salvage value amounting to $50,000.

Annuity factor of WACC 10% for 5 years = [1 - 1 / (1+0.10)5] / 0.10 = 3.7908

Present Value of Annual Cash flows = $515,000 x 3.7908 = $1,952,262

Present Value of Salvage Value = $50,000 / (1+0.10)5 = $31,046

Total Present value of cash inflows = $1,983,308

Present value of cash outflows = initial investment = $1,000,000

Net Present Value = 1,983,308 - 1,000,000 = $ 983,308